# Diffusion posterior sampling for simulation-based inference in tall data settings

### 11 Apr 2024 v1

@@ 4.2 Benchmark SBI examples
We consider three examples from the increasingly popular SBI benchmark presented in Lueckmann et al., 2021b .
@@
  • SLCP ($m=5$, $d=8$): Uniform prior and Gaussian simulator whose mean and covariance are non-linear functions of input parameter $\theta$.
@@
  • SIR ($m=2$, $d=10$): Log-Normal prior and simulator based on a set of differential equations and outputs sampled from a Binomial distribution.
@@
  • Lotka-Volterra ($m=4$, $d=20$): Log-Normal prior and simulator based on a set of differential equations and outputs sampled from a Log-Normal output.
@@
The score corresponding to each prior distribution is analytically computable; see Appendix D.555 The Log-Normal distribution can easily be transformed into a Gaussian distribution since when $\theta\sim\mathrm{LogNormal}(m,s)$ then $\log\theta\sim\mathcal{N}(m,s)$. However, contrarily to the toy models considered in Section 4.1, the analytical posterior score is not available for the chosen examples and is, therefore, learned via score-matching (considering the loss function from Section 2.1).
Note that the purpose of the experiments in this section is not to find the best score model and the best posterior approximations for each task, which has already been done by Sharrock et al., (2022) and Geffner et al., (2023), but to evaluate the robustness of the different sampling algorithms. This is why we use the same score model architecture and training hyper parameters for all tasks, which may lead to different quality for the score estimator in each example (the data dimension and shape of the posterior impact the training of the score model). The used score model is always a MLP with 3 hidden layers, trained on $N_{\mathrm{train}}$ samples over 5000 epochs using the Adam optimizer. We standardize the train data to have zero-mean and $\mathbf{I}_{m}$ covariance matrix, which ensures that all the random variables $\theta_{t}$ created during the generative process have zero-mean and $\mathbf{I}_{m}$ covariance. Further details on the model architecture and training procedure are provided in Appendix E.
@@
According to the equivalent time setting defined in Section 4.1, we use $1~{}000$ sampling steps for GAUSS and $400$ steps for JAC and LANGEVIN. We also consider clipped versions for all proposed algorithms (represented with dotted lines in all figures), where the generated samples at each step are clipped between (-3,3) so to ensure that all samples stay within the region of high probability for the standard Gaussian distribution. We introduce this numerical procedure to stabilize the JAC and LANGEVIN algorithms, that we found to be less robust than GAUSS, with quickly diverging sW for poor score estimators; see Section 4.1. However, this procedure significantly slows down the sampling procedure and can introduce some bias in the posterior approximation.
@@
Our empirical evaluation samples twenty-five [^6] We discarded outliers (max. $3\%$ of the 25 points), which correspond to a sW for GAUSS above the $99\%$ quantile (or NaNs). different ground truth parameters $\theta^{\star}\sim\lambda(\theta)$ from the prior distribution, each of which used to simulate observations $x^{\star}_{j}\sim p(x\mid\theta^{\star})$ for $j=1,\dots,n$, later plugged in as conditioning variables in the tall posterior $p(\theta\mid x^{\star}_{1:n})$. For all three examples, samples from the true tall posterior can be obtained via MCMC using the numpyro package (Phan et al.,, 2019; Bingham et al.,, 2019) and used to compute the sliced Wasserstein (sW) distance in every setting corresponding to varying $N_{\mathrm{train}}$ and $n$.
Figure 3 portrays the sW distance for each task as a function of the size $N_{\mathrm{train}}$ of the training set for the score model. Larger values of $N_{\mathrm{train}}$ are expected to yield better score estimates and thus more accurate posterior approximations. This is the case for all three examples, where we observe a decreasing tendency and convergence to 00 for all $n$. Overall, we observe that our algorithm GAUSS outperforms all others. Indeed, it scales to high $n$ values, while LANGEVIN diverges (or is non-decreasing) for $n\geq 14$ for the Lotka-Volterra and SIR examples. Note that the un-clipped version of JAC diverges in every case as soon as $n>1$, which is why we didn’t include it in the plots. GAUSS yields consistently lower distance values than LANGEVIN. JAC-clip is more or less equivalent to GAUSS, with slightly better results for SLCP. In Appendix H, Figure 11 display the same results, but as a function the number of observations $n\in[1,8,14,22,30]$ for a given score model (i.e. fixed $N_{\mathrm{train}}$). We also report results for the Maximum Mean Discrepancy (MMD) metric in Figures 12 and 13. Additionally, Figures 14 and 15 showcase the ability of the obtained posterior samples to concentrate around the ground truth parameter $\theta^{\star}$.
@@ 4.3 Inverting a non-linear model from computational neuroscience
We illustrate our proposal on a classic model from computational neuroscience and consider the Bayesian inversion of the Jansen & Rit neural mass model (JRNMM) (Jansen and Rit,, 1995). Neural mass models are non-linear models constructed based on physiologically motivated stochastic differential equations and are able to replicate oscillatory electrical signals experimentally observed with electroencephalography (EEG). Neural mass models are commonly used in large-scale simulators of the brain (Sanz Leon et al.,, 2013) and are present in several simulation studies in cognitive and clinical neuroscience (Aerts et al.,, 2018).
We follow the setup proposed by Rodrigues et al., (2021). Specifically, we consider a stochastic version of the JRNMM (Ableidinger et al.,, 2017) and use the C++ code provided by the authors of Buckwar et al., (2019). The output $x(t)$ of this generative model is a time series obtained by taking as input a set of four parameters $\theta=(C,\mu,\sigma,g)$. To help with the interpretations of the results, we provide some details on the parameters: Parameter $C$ influences the general shape of the oscillatory behavior, $(\mu,\sigma)$ drive the statistics of the signals $s(t)$ generated by the physiological model, and $g$ represents a gain factor of an amplifier (resp. attenuator) used to measure the signals and produce the actual model output $x(t)=10^{g/10}s(t)$. The coupling-effect of parameters $g$ and $(\mu,\sigma)$ on the amplitude of the output signal $x(t)$ means that the model is non-injective: the same observed signal $x^{\star}(t)$ could be generated with larger (resp. smaller) values of $g$ and smaller (resp. larger) values of $\mu$ and $\sigma$.
We now apply the different sampling algorithms GAUSS, JAC and LANGEVIN (and their clipped versions [^7] Numerical procedure used to stabilize JAC; see Section 4.2.), again with respectively 1000, 400 and 400 steps, to infer the posterior distribution of the JRNMM for a set of $n$ observations $x^{\star}_{j}\sim p(x\mid\theta^{\star})$ independently simulated with $\theta^{\star}=(C^{\star},\mu^{\star},\sigma^{\star},g^{\star})=(125,220,2000,0)$. A uniform prior distribution is placed over the range of physiologically meaningful values of the simulator parameters as done in Buckwar et al., (2019); Rodrigues et al., (2021). The posterior score is again estimated using the model architecture and training procedure from Section 4.2. We evaluate the ability of our posterior approximator to concentrate around the true parameters $\theta^{\star}$, which we quantify using the Maximum Mean Discrepancy (MMD) between the marginals of the approximate posterior and the Dirac distribution $\delta_{\theta^{\star}}$ centered at the true parameter $\theta^{\star}$.
We first consider a simplified setting for which the gain parameter is fixed $g=g^{\star}=0$ so to lift the indeterminacy on the estimation of parameters $(\mu,\sigma)$. Results are shown in Figure 4. On the left, we plot the MMD as a function of the number $n$ of observations. JAC yields values outside of the plot limits. All other methods perform as wanted, with decreasing MMD. On the right are displayed the 1D marginals corresponding to the posterior samples obtained with GAUSS for observation sets $x^{\star}_{1:n}$ of increasing size. We observe a progressive concentration around $\theta^{\star}$.
@@
Figure 4: Inference on the 3D JRNMM (fixed $g=0$). Left: MMD between the marginals of the approximate posterior obtained for for GAUSS, JAC, and LANGEVIN, and the Dirac of the true parameters $\theta^{\star}=(125,220,2000)$ (black dotted lines) used to simulate the observations $x^{\star}_{1:n}$. Right: Histograms of the 1D marginals of the posterior samples obtained with GAUSS for $30$ single observations $x^{\star}_{1},\dots,x^{\star}_{30}$ ($n=1$) and for observation sets $x^{\star}_{1:n}$ of increasing size. We see that $p(\theta\mid x^{\star}_{1:n})$ concentrates around $\theta^{\star}$.
Figure 5 displays the results for the full 4-dimensional JRNMM. On the left plot, we can see that GAUSS yields consistenty lower MMD values, with a decrease for lower $n$, LANGEVIN is unstable and JAC yields values outside of the plot limits. The 1D and 2D marginals for samples obtained for GAUSS are shown on the right. The non-decreasing behaviour of MMD for higher $n$ can be explained by the indeterminacy in the inference task for the $(\mu,\sigma)$ parameters, and is illustrated on the right plot, showing the 1D and 2D marginals for samples obtained for GAUSS. Indeed, we observe a “sharpened banana” shaped posterior distribution, containing the true parameters, concentrated around $C$ and $g$, but dispersed along the dimensions of $(\mu,\sigma)$.
Figure 5: Inference on the full JRNMM. Left: MMD between the marginals of the approximate posterior obtained for GAUSS, JAC, and LANGEVIN, and the Dirac of the true parameters ${\theta^{\star}=(125,220,2000,0)}$ used to simulate the observations $x^{\star}_{1,n}$. Right: Histograms of the 1D and 2D marginals of the posterior samples obtained with GAUSS for a single observations $x^{\star}$ ($n=1$) and for observation sets $x^{\star}_{1:n}$ of increasing $n$. We observe a progressive convergence of the inferred posterior mean towards $\theta^{\star}$ (black dots and lines), a sharpening around $(C,g)$ and a sustained dispersion along the dimensions of $(\mu,\sigma)$ due to JRNMM’s intrinsic indeterminacy.
@@ 5 Conclusion
In every experimental setting it is always desirable to leverage information from as much data as possible. Model inversions with SBI are no exception and the tall data extension considered in this paper provides a way of parsing extra observations in an efficient way. Indeed, our method only requires the training of an initial score model for a single observation to construct the score of the tall data posterior. This allows GAUSS and JAC to be more simulation efficient, avoiding the shortcomings of the augmented datasets required by NPE, NLE, and NRE. Moreover, we tackle directly the challenge of constructing a diffused version of the posterior distribution for tall data based on recently proposed second order approximations of the backward diffusion kernels. This is considerably different from what was proposed in Geffner et al., (2023), where the problem was simplified by constructing a sequence of distributions relying on the individual scores, but at the cost of drastically reducing the arsenal of samplers available for sampling from a score based generative model. Indeed, our methodology can benefit from recent advances of score-diffusion literature, such as DDIM Song et al., 2021b , to perform backward sampling of the diffusion path and obtain samples from the posterior distribution in a much faster and stable way than the annealed Langevin procedure used in Geffner et al., (2023). We have illustrated the performance of our methods on different settings of increasing complexity, starting with toy examples for which the analytic posterior and scores were known, considering three examples from the SBI benchmark, and then finally applying the methodology to a challenging non-linear model from computational neuroscience. We demonstrated the superiority of our methods in terms of sample quality for almost all examples, and the reduced computational cost and increased numerical stability in every instance. Our work has confirmed the flexibility and viability of score-diffusion methods to approximate posterior distributions conditioned on sets of variable sizes and we expect that it will encourage other researchers to explore the score-diffusion framework for SBI problems.
Acknowledgement
@@ Appendix A Proofs
since by assumption $\Lambda(\theta)$ is definite positive. This provides the following reformulation of equation (11):
$\ell_{\lambda}(\theta,x^{\star}_{1:n})=\log\int\exp\left(\sum_{j=1}^{n}\zeta_{j}(\theta)+(1-n)\zeta_{\lambda}(\theta)-\zeta_{\rm{all}}(\theta)\right)\mathcal{N}(\theta_{0};\Lambda^{-1}(\theta)\eta(\theta),\Lambda^{-1}(\theta))\mathrm{d}\theta_{0}\,,$
which concludes the proof.
@@
Let $\zeta(\mu,\Sigma^{-1})=-\left(m\log 2\pi-\log|\Sigma^{-1}|+\mu^{\top}\Sigma^{-1}\mu\right)/2$. Then,
$\nabla_{\theta}\zeta_{j}(\theta,x_{j})$ $=\nabla_{\theta}\zeta(\mathrm{\mu}_{t}(\theta,x_{j}),\Sigma^{-1}_{t}(\theta,x_{j}))$
$=\nabla_{\mu}\zeta(\mathrm{\mu}_{t}(\theta,x_{j}),\Sigma^{-1}_{t}(\theta,x_{j}))^{\top}\nabla_{\theta}\mu_{t}(\theta,x_{j})$
$\hskip 142.26378pt+\nabla_{\Sigma^{-1}}\zeta(\mathrm{\mu}_{t}(\theta,x_{j}),\Sigma^{-1}_{t}(\theta,x_{j}))\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})\,,$
@@
$\nabla_{\mu}\zeta(\mu,\Sigma^{-1})$ $=-\Sigma^{-1}\mu$
$\nabla_{\Sigma}^{-1}\zeta(\mu,\Sigma^{-1})$ $=(1/2)\left(\Sigma-\mu\mu^{\top}\right)\,.$
@@
$\nabla_{\theta}\zeta_{j}(\theta,x_{j})$ $=-\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\Sigma^{-1}_{t}(\theta,x_{j})^{\top}\nabla_{\theta}\mathrm{\mu}_{t}(\theta,x_{j})$
$\hskip 56.9055pt+(1/2)\left(\Sigma_{t}(\theta,x_{j})-\mathrm{\mu}_{t}(\theta,x_{j})\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})\,.$
Note that $\nabla_{\theta}\mathrm{\mu}_{t}(\theta,x_{j})=(\sqrt{\alpha_{t}}/\upsilon_{t})\Sigma_{t}(\theta,x_{j})$, which leads to
$\nabla_{\theta}\zeta_{j}(\theta,x_{j})$ $=-\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\mathrm{\mu}_{t}(\theta,x_{j})+(1/2)\left(\Sigma_{t}(\theta,x_{j})-\mathrm{\mu}_{t}(\theta,x_{j})\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})$
$=-\upsilon_{t}^{-1}\theta-\nabla_{\theta}\log p_{t}(\theta){x_{j}}+(1/2)\left(\Sigma_{t}(\theta,x_{j})-\mathrm{\mu}_{t}(\theta,x_{j})\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})\,.$
@@
$\hskip 28.45274pt+(1/2)\left[\sum_{j=1}^{n}\left(\Sigma_{t}(\theta,x_{j})-\mathrm{\mu}_{t}(\theta,x_{j})\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})\right]$
@@
$\hskip 28.45274pt+(1/2)\left[\sum_{j=1}^{n}\left(\Sigma_{t}(\theta,x_{j})-\mathrm{\mu}_{t}(\theta,x_{j})\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})\right]$
@@
$\nabla_{\theta}\eta(\theta)=\sum_{j=1}^{n}\Sigma^{-1}_{t}(\theta,x_{j})\nabla_{\theta}\mathrm{\mu}_{t}(\theta,x_{j})+\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})$
@@
$=\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\mathbf{I}_{m}+\sum_{j=1}^{n}\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})+(1-n)\mathrm{\mu}_{t,\lambda}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,\lambda}(\theta)\,,$
@@
$\nabla_{\theta}\eta(\theta)\eta(\theta)=\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\eta(\theta)+\left[\sum_{j=1}^{n}\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})+(1-n)\mathrm{\mu}_{t,\lambda}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,\lambda}(\theta)\right]\eta(\theta)\,.$ (13)
@@
$\sqrt{\alpha_{t}}\eta(\theta)$ $=\sum_{j=1}^{n}\Sigma^{-1}_{t}(\theta,x_{j})\left(\theta+\upsilon_{t}\nabla_{\theta}\log p_{t}(\theta){x_{j}}\right)+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\left(\theta+\upsilon_{t}\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}\right)$
$=\Lambda(\theta)\theta+\upsilon_{t}\left[\sum_{j=1}^{n}\Sigma^{-1}_{t}(\theta,x_{j})\nabla_{\theta}\log p_{t}(\theta){x_{j}}+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}\right]\,,$
@@
$\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{1:n})=\Lambda(\theta)^{-1}\left[\sum_{j=1}^{n}\Sigma^{-1}_{t}(\theta,x_{j})\nabla_{\theta}\log p_{t}(\theta){x_{j}}+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}\right]+F(\theta,x^{\star}_{1:n})\,,$
@@
$+(1/2)\left[\sum_{j=1}^{n}\left(\Sigma_{t}(\theta,x_{j})-\mathrm{\mu}_{t}(\theta,x_{j})\mathrm{\mu}_{t}(\theta,x_{j})^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t}(\theta,x_{j})\right]$

### 07 Jun 2024 v2

@@ 4.2 Benchmark SBI examples
We consider three examples from the popular SBI benchmark presented in Lueckmann et al., (2021):
@@
  • SLCP ($m=5$, $d=8$): Uniform prior and Gaussian simulator, whose mean and covariance are non-linear functions of the input parameters $\theta$.
@@
  • SIR ($m=2$, $d=10$): Log-Normal prior and simulator based on a set of differential equations that outputs sampled from a Binomial distribution.
@@
  • Lotka-Volterra ($m=4$, $d=20$): Log-Normal prior and simulator based on a set of differential equations that outputs sampled from a Log-Normal distribution.
@@
The score corresponding to each prior distribution is analytically computable; see Appendix D. However, contrarily to the toy models considered in Section 4.1, the analytical posterior score is not available for the chosen examples and is, therefore, learned via score-matching as explained in Section 2. The goal of this experiment is to evaluate the robustness of the different sampling algorithms to this learning setting.
@@
The used score model is always a MLP with 3 hidden layers, trained on $N_{\mathrm{train}}$ samples over 5000 epochs using the Adam optimizer. Further details are provided in Appendix E. According to the equivalent time setting defined in Section 4.1, we use $1~{}000$ sampling steps for GAUSS and $400$ steps for JAC and LANGEVIN. We also consider clipped versions for all proposed algorithms (represented with dashed lines), where we ensure that the samples generated at every step stay within the region of high probability for the standard Gaussian distribution. We introduce this numerical procedure to stabilize the JAC and LANGEVIN algorithms, that we found to be less robust than GAUSS, with quickly diverging sW for poor score estimators; see Section 4.1. Note that this procedure significantly slows down the sampling procedure and can introduce some bias in the posterior approximation.
@@
Our empirical evaluation samples twenty-five [^4] We removed outliers, which correspond to a sW for GAUSS above the $99\%$ quantile (or NaN values). different ground truth parameters $\theta^{\star}\sim\lambda(\theta)$ from the prior distribution, each of which used to simulate observations $x^{\star}_{j}\sim p(x\mid\theta^{\star})$ for $j=1,\dots,n$, later plugged in as conditioning variables in the tall posterior $p(\theta\mid x^{\star}_{1:n})$. For all three examples, samples from the true tall posterior can be obtained via MCMC using the numpyro package (Phan et al.,, 2019; Bingham et al.,, 2019) and used to compute the sliced Wasserstein (sW) distance in every setting corresponding to varying $N_{\mathrm{train}}$ and $n$. In Appendix H, we also report results for the Maximum Mean Discrepancy (MMD) and the Classifier-Two-Sample Test (C2ST) metrics from the sbibm package.
@@
Figure 3 portrays the sW distance for each task as a function of the size $N_{\mathrm{train}}$ of the training set for the score model. Larger values of $N_{\mathrm{train}}$ are expected to yield better score estimates and thus more accurate posterior approximations. This is represented by a decreasing tendency towards 00 of the sW. Overall, we observe that GAUSS outperforms all other algorithms. It yields consistently lower distance values than LANGEVIN and scales to high $n$ values, while LANGEVIN diverges for $n\geq 14$ for the Lotka-Volterra and SIR examples. Note that JAC diverges in every case as soon as $n>1$, which is why we didn’t include it in the plots. On the other hand, JAC-clip is more or less equivalent to GAUSS, with slightly better results for SLCP. We refer the reader to Appendix I for results on additional tasks from the SBI benchmark. Here, all samplers fail to accurately infer the tall posterior in more complex data settings, which opens up the question on the scalability the Gaussian approximation. One limitation consists in the multiple score network evaluations, which can lead to an accumulation of errors with increasing $n$ (see Appendix H). As a response, Appendix L.2 includes results of a partially factorized version of our method, as proposed by Geffner et al., (2023), allowing the best trade-off between simulation cost and number of network evaluations.
@@ 4.3 Inverting a non-linear model from computational neuroscience
We illustrate our proposal on a classic model from computational neuroscience and consider the Bayesian inversion of the Jansen & Rit neural mass model (JRNMM) (Jansen and Rit,, 1995). Neural mass models are non-linear models constructed based on physiologically motivated stochastic differential equations and are able to replicate oscillatory electrical signals experimentally observed with electroencephalography (EEG) and are commonly used in large-scale simulators of the brain (Sanz Leon et al.,, 2013). We follow the setup proposed by Rodrigues et al., (2021). Specifically, we consider a stochastic version of the JRNMM (Ableidinger et al.,, 2017) with a uniform prior placed over the range of physiologically meaningful values of the simulator parameters and use the code provided by Buckwar et al., (2019). The output $x(t)$ of this generative model is a time series obtained by taking as input a set of four parameters $\theta=(C,\mu,\sigma,g)$. All parameters have a physiological interpretation and we invite the reader to check the above references for a detailed explanation on each one of them. It is worth mentioning, however, that there exists a coupling-effect of parameters $g$ and $(\mu,\sigma)$ on the amplitude of the output signal $x(t)$, meaning that the model is non-identifiable: the same observed signal $x^{\star}(t)$ could be generated with larger (smaller) values of $g$ and smaller (larger) values of $\mu$ and $\sigma$. This indeterminacy makes the posterior inference problem ill-posed and we show next how a tall data posterior can give more precise information on how to invert the simulator.
[…]
We illustrate the ability of our tall posterior to concentrate around the true parameters $\theta^{\star}$ used to simulate the observations $x^{\star}_{1:n}$ as $n$ increases. This is quantified with the MMD between the posterior marginals and the Dirac distribution $\delta_{\theta^{\star}}$, computed on 10 0001000010\,00010 000 samples obtained with GAUSS. Figure 4 portrays the results for a simplified setting in which the gain parameter is fixed at $g=0$ so to lift the indeterminacy on $(\mu,\sigma)$. We observe as expected a progressive concentration around $\theta^{\star}$, with sharper posterior marginals and decreasing MMD. In Appendix J, we include the results obtained for the full (4D) JRNMM for which the tall posterior takes a ”sharpened banana” shape: while we observe a concentration around $(C,g)$, the dispersion along the dimensions of $(\mu,\sigma)$ is sustained.
@@
Figure 4: Inference on the 3D JRNMM (fixed $g=0$) with GAUSS. Left: MMD between the marginals of the approximate posterior and the Dirac of the true parameters $\theta^{\star}$ (black dashed lines) used to simulate the observations $x^{\star}_{1:n}$. Right: Histograms of the 1D marginals of the inferred posterior samples for $30$ single observations $x^{\star}_{1},\dots,x^{\star}_{30}$ ($n=1$) and observation sets $x^{\star}_{1:n}$ of increasing size.
@@ 5 Conclusion
The tall data setting considered in this paper provides a way of parsing extra observations in an efficient way for SBI. Our method only requires the training of an initial score model amortized for single observations to construct the score of the tall data posterior. This allows GAUSS and JAC to be more simulation efficient, avoiding the shortcomings of the augmented datasets required by NPE. Moreover, we tackle directly the challenge of constructing a diffused version of the posterior distribution for tall data based on a second order approximations of the backward diffusion kernels. Further research will investigate the scalability of this approximation both theoretically and practically. Our method benefits from recent advances of score-diffusion literature, such as DDIM (Song et al., 2021b, ), to perform backward sampling of the diffusion path and obtain samples from the posterior distribution in a much faster and stable way than the annealed Langevin procedure used in Geffner et al., (2023). We have illustrated the performance of our methods on different settings of increasing complexity, considering in particular three examples from the SBI benchmark, and applying our methodology to a challenging non-linear model from computational neuroscience. We demonstrated the superiority of our methods in terms of sample quality for almost all examples, and the reduced computational cost and increased numerical stability in every instance. We expect that our work will encourage other researchers to explore the score-diffusion framework for SBI problems.
Acknowledgements
Julia Linhart is recipient of the Pierre-Aguilar Scholarship and thankful for the funding of the Capital Fund Management (CFM). The work of G. Cardoso was supported by the Investment of the Future Grant, ANR-10-IAHU-04, from the government of France through the Agence National de la Recherche. Alexandre Gramfort thanks the support of the ANR BrAIN (ANR-20-CHIA0016) grant.
@@ Appendix A Proofs
since by assumption $\Lambda(\theta)$ is definite positive. This provides the following reformulation of equation (10), which concludes the proof:
$\ell_{\lambda}(\theta,x^{\star}_{1:n})=\log\int\exp\left(\sum_{j=1}^{n}\zeta_{j}(\theta)+(1-n)\zeta_{\lambda}(\theta)-\zeta_{\rm{all}}(\theta)\right)\mathcal{N}(\theta_{0};\Lambda^{-1}(\theta)\eta(\theta),\Lambda^{-1}(\theta))\mathrm{d}\theta_{0}\,.$
@@
Let $\zeta(\mu,\Sigma^{-1})=-\left(m\log 2\pi-\log|\Sigma^{-1}|+\mu^{\top}\Sigma^{-1}\mu\right)/2$. Then
$\nabla_{\theta}\zeta_{j}(\theta,x_{j})$ $=\nabla_{\theta}\zeta(\mathrm{\mu}_{t,j}(\theta),\Sigma^{-1}_{t,j}(\theta))$
$=\nabla_{\mu}\zeta(\mathrm{\mu}_{t,j}(\theta),\Sigma^{-1}_{t,j}(\theta))^{\top}\nabla_{\theta}\mu_{t,j}(\theta)+\nabla_{\Sigma^{-1}}\zeta(\mathrm{\mu}_{t,j}(\theta),\Sigma^{-1}_{t,j}(\theta))\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)\,,$
@@
$\nabla_{\mu}\zeta(\mu,\Sigma^{-1})=-\Sigma^{-1}\mu\quad\text{and}\quad\nabla_{\Sigma}^{-1}\zeta(\mu,\Sigma^{-1})=(1/2)\left(\Sigma-\mu\mu^{\top}\right)\,.$
@@
$\nabla_{\theta}\zeta_{j}(\theta,x_{j})$ $=-\mathrm{\mu}_{t,j}(\theta)^{\top}\Sigma^{-1}_{t,j}(\theta)^{\top}\nabla_{\theta}\mathrm{\mu}_{t,j}(\theta)$
$\hskip 56.9055pt+(1/2)\left(\Sigma_{t,j}(\theta)-\mathrm{\mu}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)\,.$
Note that $\nabla_{\theta}\mathrm{\mu}_{t,j}(\theta)=(\sqrt{\alpha_{t}}/\upsilon_{t})\Sigma_{t,j}(\theta)$, which leads to
$\nabla_{\theta}\zeta_{j}(\theta,x_{j})$ $=-\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\mathrm{\mu}_{t,j}(\theta)+(1/2)\left(\Sigma_{t,j}(\theta)-\mathrm{\mu}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)$
$=-\upsilon_{t}^{-1}\theta-\nabla_{\theta}\log p_{t}(\theta){x_{j}}+(1/2)\left(\Sigma_{t,j}(\theta)-\mathrm{\mu}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)\,.$
@@
$\hskip 28.45274pt+(1/2)\left[\sum_{j=1}^{n}\left(\Sigma_{t,j}(\theta)-\mathrm{\mu}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)\right]$
@@
$\hskip 28.45274pt+(1/2)\left[\sum_{j=1}^{n}\left(\Sigma_{t,j}(\theta)-\mathrm{\mu}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)\right]$
@@
$\nabla_{\theta}\eta(\theta)=\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\nabla_{\theta}\mathrm{\mu}_{t,j}(\theta)+\mathrm{\mu}_{t,j}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)$
@@
$=\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\mathbf{I}_{m}+\sum_{j=1}^{n}\mathrm{\mu}_{t,j}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)+(1-n)\mathrm{\mu}_{t,\lambda}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,\lambda}(\theta)\,,$
@@
$\nabla_{\theta}\eta(\theta)\eta(\theta)=\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\eta(\theta)+\left[\sum_{j=1}^{n}\mathrm{\mu}_{t,j}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)+(1-n)\mathrm{\mu}_{t,\lambda}(\theta)^{\top}\nabla_{\theta}\Sigma^{-1}_{t,\lambda}(\theta)\right]\eta(\theta)\,.$ (13)
@@
$\sqrt{\alpha_{t}}\eta(\theta)$ $=\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\left(\theta+\upsilon_{t}\nabla_{\theta}\log p_{t}(\theta){x_{j}}\right)+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\left(\theta+\upsilon_{t}\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}\right)$
$=\Lambda(\theta)\theta+\upsilon_{t}\left[\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\nabla_{\theta}\log p_{t}(\theta){x_{j}}+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}\right]\,,$
@@
$\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{1:n})=\Lambda(\theta)^{-1}\left[\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\nabla_{\theta}\log p_{t}(\theta){x_{j}}+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}\right]+F(\theta,x^{\star}_{1:n})\,,$
@@
$+(1/2)\left[\sum_{j=1}^{n}\left(\Sigma_{t,j}(\theta)-\mathrm{\mu}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)^{\top}\right)\nabla_{\theta}\Sigma^{-1}_{t,j}(\theta)\right]$